OpenStudy (anonymous):

Enrollment at a local university is 8500. There are one third as many students over the age 25 as there are 25 and under. how many students are 25 and under and how many are over over 25?

OpenStudy (zarkon):

6375,2125

OpenStudy (anonymous):

let over 25 = a and under 25 =b a=1/3b 3a=b also a+b=8500 a+3a=8500 4a = 8500 a = 2,125

OpenStudy (anonymous):

That dont seem right Zarkon

OpenStudy (anonymous):

so b=6375

OpenStudy (zarkon):

looks good to me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@ soccerdoggie, where do the b=6375 what do you after thar

OpenStudy (anonymous):

where do the b=6375 come in at?

OpenStudy (zarkon):

we gave you both answers: 6375 that are 25 and under 2125 that are over 25